Night report 2021-08-23

Observers: Priscila Pessi

Support: Claudia Gutiérrez

TAT: Michael Fulton, Takashi Nagao, Jacob Teffs

EFOSC Calibrations:

Bias, domeflats

SOFI Calibrations:

Weather:

Start: humidity ~26%, wind ~11.4m/s, seeing ~1.4"

Middle: humidity ~%, wind ~m/s, seeing ~"

End: humidity ~%, wind ~m/s, seeing ~"

Comments (times in UT):

At around 00:50UT cirrus made an appearence

At 02:45UT we closed the telescope due to high winds

At 04:17UT we open again. We try to do image analysis but the system fails. After starting with a standard we hadto close again due to clouds.